с) х - 3x° — 4х + 12
Answer #1

let y = 23-30% - 4n+12 = mº (-3) - 4 (m3) = (0%-4) (n-3) = 81+a)02-a)(x-3) 3 - (+2)(x-3)(x -3) zeros of y are 175-22.31